# Modeling Cell Dynamics and Interactions with Unbalanced Mean Field Schrödinger Bridge

### v2

@@ C.2 Hyperparameters Selection and Loss Weighting
$\lambda_{m}$ 1 1 1 1 1
$\lambda_{d}$ 1 1 1 1 1
$\lambda_{r}$ $1\times 10^{3}$ $1\times 10^{3}$ $1\times 10^{3}$ $1\times 10^{3}$ $1\times 10^{3}$
$\lambda_{f}$ $1\times 10^{4}$ $1\times 10^{4}$ $1\times 10^{4}$ $1\times 10^{4}$ $1\times 10^{4}$
$\lambda_{w}$ 10 10 10 10 10
$d_{\text{cutoff}}$ 0.5 300 100 100 2
Epochs 200 1000 500 500 200
C.3 Visualization
@@ Appendix D Mathematical Details
From Definition 4.1 we obtain
@@
The initial positions $\mathbf{X}_{0}^{i}$ are independently and identically distributed (i.i.d.) with a common density $\rho_{0}(\mathbf{x})\in L^{1}(\mathbb{R}^{d})\cap L^{\infty}(\mathbb{R}^{d})$, and the initial weights are set as $w_{i}(0)=1$ for all $i=1,\dots,N$. The functions $\mathbf{b}$, $g$, $\phi$, and $\Phi$ are Lipschitz continuous and bounded: specifically, $\mathbf{b}$ and $g$ are Lipschitz in $\mathbf{x}$ uniformly in $t$, $k(\mathbf{x},\mathbf{y})$ is Lipschitz in both arguments and bounded, and $\nabla_{x}\Phi$ is Lipschitz continuous and bounded. The diffusion coefficient $\sigma(t)$ is continuous and bounded on $[0,T]$. The empirical measure $\mu_{t}^{N}=\frac{1}{N}\sum_{i=1}^{N}w_{i}(t)\delta_{\mathbf{X}_{t}^{i}}$ converges to a deterministic limit $\rho(\mathbf{x},t)$. The system in Proposition 5.1 satisfies argument as $N\to\infty$, i.e., $\frac{1}{N-1}\sum_{j\neq i}w_{j}(t)f(\mathbf{X}_{t}^{j},t)\rightarrow\int_{\mathbb{R}^{d}}f(\mathbf{x},t)\rho(\mathbf{x},t)$, where $f$ is an arbitrary test function.
@@
The argument we assume here is related to the notions of “chaos” and “propagation of chaos” and it has a rich theory in mathematics (Jabin & Wang, 2017; Chaintron & Diez, 2021). To rigorously prove such an argument, we refer to some related works (Fournier et al., 2014; Feng & Wang, 2024; Duteil, 2022; Ben-Porat et al., 2024; Ayi & Duteil, 2021).
@@
Appendix E Limitations and Further Directions
While CytoBridge offers valuable insights into incorporating cell-cell interaction with unbalanced stochastic dynamics, several aspects could benefit from further exploration. Firstly, CytoBridge minimizes the upper bound of the energy term in the UMFSB problem. Directly optimizing the original UMFSB formulation remains an interesting question. Potential solutions may involve leveraging neural SDEs or the integration-by-parts strategy proposed by (Zhang et al., 2025a) as viable approaches. Secondly, the current neural network parameterization of cellular interaction terms is based on techniques used to reduce degrees of freedom, such as RBF expansion. Future work could explore sparse representation methods to replace it for improved expressive power. Thirdly, the training of CytoBridge involves multiple stages. Simplifying the training process by incorporating optimality conditions (e.g., HJB equations) presents a promising research direction. Finally, extending the concept of flow matching to this context and developing simulation-free training methods for stochastic, growth/death, and interaction dynamics could also advance the CytoBridge’s applicability.

### v3

@@ C.2 Hyperparameters Selection and Loss Weighting
$\lambda_{m}$ 1 1 1 1 1 1
$\lambda_{d}$ 1 1 1 1 1 1
$\lambda_{r}$ $\times 10^{3} 1×1$$\times 10^{3}$ $\times 10^{3} 1×1$$\times 10^{3}$ $\times 10^{3} 1×1$$\times 10^{3}$
$\lambda_{f}$ $\times 10^{4} 1×1$$\times 10^{4}$ $\times 10^{4} 1×1$$\times 10^{4}$ $\times 10^{4} 1×1$$\times 10^{4}$
$\lambda_{w}$ 10 10 10 10 10 10
$d_{\text{cutoff}}$ 0.5 300 100 100 2 0.3
Epochs 200 1000 500 500 200 500
C.3 Scalability and Computational Efficiency
We conducted an evaluation of the scalability and computational efficiency of CytoBridge on the embryoid body data by extending the input from 50 to 100 PCs. As shown in Table˜20, CytoBridge achieves the lowest $\mathcal{W}_{1}$ distance across all time points, demonstrating that CytoBridge remains effective at higher dimensionality. Regarding the computational efficiency, CytoBridge maintains comparable to other neural ODE-based methods with a training time of 11 minutes and a peak GPU memory usage of 6.3 GB.
Table 20: Performance on 100D embryoid body data. We show the Wasserstein distance ($\mathcal{W}_{1}$) at four time points, alongside total runtime and peak memory usage. Bold indicates the best result.
$t=1$ $t=2$ $t=3$ $t=4$ Time Memory (GB)
Model $\mathcal{W}_{1}$ $\mathcal{W}_{1}$ $\mathcal{W}_{1}$ $\mathcal{W}_{1}$
SF2M (Tong et al., 2024b) 11.333 12.982 13.718 14.945 4min 43s 0.7
Meta FM (Atanackovic et al., 2025) 11.699 13.398 14.037 14.727 5min 50s 2.2
MMFM (Rohbeck et al., 2025) 13.150 14.135 14.441 14.907 4min 38s 0.6
Metric FM (Kapusniak et al., 2024) 10.806 12.348 13.622 16.801 2min 16s 4.2
UOT-FM (Eyring et al., 2024) 10.757 12.799 13.761 15.657 3min 27s 0.6
uAM (Neklyudov et al., 2023) 13.628 18.315 20.309 22.973 38s 1.5
MIOflow (Huguet et al., 2022) 11.387 12.331 11.905 12.908 6min 19s 0.7
TIGON (Sha et al., 2024) 10.547 12.926 13.897 14.535 7min 33s 0.7
DeepRUOT (Zhang et al., 2025a) 10.226 11.110 11.544 12.424 10min 17s 3.7
CytoBridge (Ours) 10.217 11.070 11.505 12.368 11min 26s 6.3
C.4 Visualization
@@ Appendix D Mathematical Details
From Definition˜4.1 we obtain
@@
The initial positions $\mathbf{X}_{0}^{i}$ are independently and identically distributed (i.i.d.) with a common density $\rho_{0}(\mathbf{x})\in L^{1}(\mathbb{R}^{d})\cap L^{\infty}(\mathbb{R}^{d})$, and the initial weights are set as $w_{i}(0)=1$ for all $i=1,\dots,N$. The functions $\mathbf{b}$, $g$, $\phi$, and $\Phi$ are Lipschitz continuous and bounded: specifically, $\mathbf{b}$ and $g$ are Lipschitz in $\mathbf{x}$ uniformly in $t$, $k(\mathbf{x},\mathbf{y})$ is Lipschitz in both arguments and bounded, and $\nabla_{x}\Phi$ is Lipschitz continuous and bounded. The diffusion coefficient $\sigma(t)$ is continuous and bounded on $[0,T]$. The empirical measure $\mu_{t}^{N}=\frac{1}{N}\sum_{i=1}^{N}w_{i}(t)\delta_{\mathbf{X}_{t}^{i}}$ converges to a deterministic limit $\rho(\mathbf{x},t)$. The system in Proposition˜5.1 satisfies argument as $N\to\infty$, i.e., $\frac{1}{N-1}\sum_{j\neq i}w_{j}(t)f(\mathbf{X}_{t}^{j},t)\rightarrow\int_{\mathbb{R}^{d}}f(\mathbf{x},t)\rho(\mathbf{x},t)$, where $f$ is an arbitrary test function.
@@
The argument we assume here is related to the notions of “chaos” and “propagation of chaos” and it has a rich theory in mathematics (Jabin and Wang, 2017; Chaintron and Diez, 2021). To rigorously prove such an argument, we refer to some related works (Fournier et al., 2014; Feng and Wang, 2024; Duteil, 2022; Ben-Porat et al., 2024; Ayi and Duteil, 2021).
@@
Appendix E More Background on Trajectory Inference
In this section, we provide more background on the trajectory inference task. In single-cell transcriptomics, methods vary depending on the data type:
Single Time-Point Snapshot Data: In this context, trajectory inference methods are broadly categorized into two groups. The first type is pseudotime-based methods (Trapnell et al., 2017; Street et al., 2018; Wolf et al., 2019), which infer trajectories by ordering cells along a pseudotime axis but often require specifying a starting point, limiting their flexibility. The second type is RNA velocity-based methods (Bergen et al., 2020; Qiao and Huang, 2021; Gayoso et al., 2024), which leverage spliced and unspliced counts to estimate dynamics but are constrained by the need for such data. Both approaches infer dynamics from existing data without generating new cell states.
Multi Time-Point Snapshot Data: Our work addresses a distinct scenario involving single-cell sequencing data across multiple time points. Due to the destructive nature of technology, this can be framed as a generative modeling task, where dynamics are inferred from distributions at different time points and could be interpolated at unseen time points. Once learned, these dynamics enable the generation of intermediate cell states. Methods like optimal transport, which have gained significant attention in computational systems biology and machine learning, are well-suited for this task (Zhang et al., 2025c).
Regarding evaluation and comparability, our generative approach, unlike pseudotime or RNA velocity methods, evaluates performance using metrics for generated distributions, which traditional methods cannot produce. This fundamental difference in data and modeling objectives, along with our hold-out experiments, renders direct comparisons with traditional trajectory inference methods infeasible. (Zhang et al., 2025b) discusses modeling approaches for different data types. We summarizes these distinctions in Table˜21.
Table 21: Comparison of Trajectory Inference Methodologies
Method Data Key Generative Evaluation
Pseudotime Single snapshot or merged snapshots without using time labels Infers trajectories by ordering cells; requires starting point No Pseudotime accuracy with true time labels
RNA Velocity Single snapshots with spliced/unspliced counts Estimates dynamics using RNA splicing kinetics No Velocity consistency
Optimal Transport Multi-time-point snapshots Infers dynamics and generates new cell states Yes Generated distribution metrics